Y= c+I+g


C= 900+0.5(y-t)


I= 750-30r


T=800


G=1200


Md=Ms


Ms= 1500


Mt=0.7y


MSP=-80r


Derive ( Md/P) the demand for real money balances equation

1
Expert's answer
2022-03-15T12:11:14-0400

Solution

"Y = C + I + G + X - M"

"Y=900+0.5(Y-T)+750-30r+1200"

"Y=900+0.5Y-0.5T+750-30r+1200"

"0.5Y=2850-30r"

"Y=\frac{2850-30r}{0.5}"